OpenStudy (anonymous):

If h(x)=6-x what is the value of (h*h)(10)?

OpenStudy (kainui):

It's probably better to just consider that (h*h)(10) really just means h(h(10)).

OpenStudy (anonymous):

Choices are A. -4 B. -2 C. 10 D. 16

OpenStudy (trojanpoem):

(h*h)(x) = (6-x)^2 (h*h)(10) = (-4)^2 = 16

OpenStudy (anonymous):

Thank you so much. The way you set it up helped me understand the problem better